OpenStudy (anonymous):

use elimination method to solve system of equation a+8b=17 and a+3b=2 would it be no solutions, many solutions or one solution i must give ordered pair Please help thanks

OpenStudy (anonymous):

would it look like this? a+8b=17 -(a+3b=2)= -a-3b+-2 you then subtract a+8b=17 -a-3b=-2 is that right so far?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

would it look like this? a+8b=17 -(a+3b=2)= -a-3b+-2 you then subtract a+8b=17 -a-3b=-2 is that right so far @Hero

OpenStudy (anonymous):

is this the answer (11,-3) @Hero

OpenStudy (anonymous):

@JamesJ can you please help me?

OpenStudy (whpalmer4):

a + 8b = 17 - (a + 3b = 2) ----------- 0a + 5b = 15 5b = 15, solve for b. Then use the value of b in one of the equations to find a. (11,-3) is not correct.

OpenStudy (anonymous):

is the answer (7,3) and thank you so much for your help @whpalmer4

OpenStudy (whpalmer4):

You're half right. Check your answer in the first equation: a+8b = 17 7+8(3) = 17?

OpenStudy (whpalmer4):

show your work for finding a, maybe we can spot where you went wrong...

OpenStudy (anonymous):

were did you get the 7 from? @whpalmer4

OpenStudy (anonymous):

a+8(3)=17 a+24=17 a-24=-24 ---------- a=-7 thats what i got @whpalmer4

OpenStudy (anonymous):

@whpalmer4 is that the right answer? (-7,3)

OpenStudy (whpalmer4):

Let's check it: -7 + 8(3) = 17 -7 + 3(3) = 2 both solutions work in both equations, so the answer is correct! good job!